# A Non-Parametric Sensitivity Analysis for Bounding Bias in Hybrid Control Trials

### v1

@@ 2.2.1 Efficient Estimation
We rely on an efficient estimator $\hat{\Psi}$ for the trial-specific average treatment effect proposed by Li et al. [28]. Although others exist, we focus on this estimator because it is an efficient estimator. This property aligns with our goal to alleviate trial constraints since efficient estimators permit smaller sample sizes while preserving standard methods of inference and robust behavior across data-generating processes [25]. This estimator is also doubly robust, making it less prone to model misspecification bias.
The efficient influence function for this estimand (as derived by Li et al. [28]) is as follows:
@@
When one assumes mean exchangeability and has a binarized outcome $Y$, $r(X)$ reduces to 1. In this paper, we will restrict the outcome to be binary so that $r(X)$ does not require estimation and the results can be presented more economically. However, the results should still extend well to other cases (e.g. continuous or even censored outcomes). For example, the value of $r(X)$ only affects efficiency and not unbiased estimation, and $r(X)$ can be estimated well in a scenario where a continuous outcome is of interest as shown in Li et al. [28].
The estimator given above is a typical efficient estimating equations (EEE, or “AIPW-style”) estimator. Although we use EEE here for accessibility and consistency with the double machine learning literature, other (asymptotically equivalent) approaches for efficient estimation such as targeted maximum likelihood estimation (TMLE) can also be used [25].
@@ 3 Methods
The issue at hand is that this estimator, like others, can only target the observable estimand which relies on the strict mean exchangeability assumption in order to get any benefit from the external control subjects. Although there are approaches such as the test-then-pool method to detect violations of this assumption, they are typically simple and cannot recognize every violation [42]. Further, preliminary tests may identify a violation of the mean exchangeability assumption, but it’s also possible that the violations only lead to small amounts of bias. In these cases, the external data may have been useful, providing greater benefits than costs, but not used due to the black-and-white nature of significance tests. More complicated methods of dynamic borrowing include Bayesian methods, but simulation studies have shown that power gains are limited when strict type-one error rates are enforced [24]. Due to the inevitability and uncertain quantification of this bias, hybrid control trials are typically discouraged by major regulatory bodies [1].
@@ 3.1 Omitted Variable Bias
This idea generalizes to any hybrid control trial and recontextualizes the issue to be an omitted variable bias problem. In particular, $\Psi$ no longer identifies $\Psi^{*}$ since it does not include the relevant covariates. Instead, $\tilde{\Psi}$, denoting the statistical estimand that conditions on $\tilde{X}$, identifies the true causal parameter $\Psi^{*}$. This discrepancy introduces the bias $|\Psi^{*}-\Psi|=|\tilde{\Psi}-\Psi|$. By bounding this bias, we can estimate how far $\Psi$ is from $\tilde{\Psi}$ when unobserved confounding is present. In the rest of this paper, we also use the tilde to denote long regressions (e.g. $\tilde{\mu}(A,\tilde{X})=E[Y|\tilde{X},A=a,D=1]$) in which the full set of confounders $\tilde{X}$ is used instead of the short regressions in which only the observed covariates $X$ are used.
In what follows, our goal will be to find an upper-bound $|\Psi^{*}-\Psi|\leq B$, quantifying the maximum bias that could be introduced if the mean exchangeability assumption were violated. We will then derive an estimator $\hat{B}$ for $B$, the bias bound.
@@ 3.2 The Bias Bound $B$
We leverage prior omitted variable bias methodologies for developing sensitivity analyses based on the Riesz representation of the causal estimand [14]. Prior sensitivity analyses have been developed for a variety of estimands in observational studies such as the average treatment effect and weighted average incremental effects. In particular, this method works for a class estimands that can be written as a linear functional of the conditional expectation function of the outcome, also known as the outcome regression. These estimands can be written in the form: $\Psi(P_{0})=E[m(\mu)(O)]$ where $\mu$ is the outcome regression based on $P_{0}$ and $m$ is some linear functional.
@@ 3.2.1 Riesz Representation
The Riesz representation theorem states that within a Hilbert space any continuous linear function that maps to a real number can be written as the inner product of the argument to the function and some other element in the Hilbert space [35, 26]. Using this theorem, we can then rewrite estimands of the form given above as $E[m(\mu)(O)]=E[\mu(O)\alpha(O)]$ where $\alpha$ is defined as the Riesz representer.
As an example, we can consider the well-known estimand for the average treatment effect of a binary treatment: $\Psi(P_{0})=E[E[Y|A=1,X]-E[Y|A=0,X]]$. The ATE can be defined as a linear functional of the outcome regression with $m(\mu)(O)=\mu(1,X)-\mu(0,X)$. Standard conditioning arguments show that the Riesz representer for the ATE is $\frac{A}{P(A=1|X)}-\frac{1-A}{P(A=0|X)}$, the usual inverse propensity weights. In this way, the Riesz representation motivates the use of such weighted estimators [26].
Chernozhukov et al. [14] uses this representation to define a bias bound $B$ for any Riesz-representable estimand, dependent on the residuals between the long and short formulations of the outcome regression and Riesz representer. Since the short regressions are projections of the long regressions onto a smaller space of functions depending only on $X$, these short regressions are orthogonal to the additional regression components living outside of this smaller subspace. Relying on the orthogonality of $\mu$ and $\alpha$ on these residuals, it is shown that the bias (in terms of the difference of the long and short estimands) is bounded: $\Psi-\tilde{\Psi}=E[(\mu(O)-\tilde{\mu}(O))(\alpha(O)-\tilde{\alpha}(O))]\leq\|\mu-\tilde{\mu}\|\|\alpha-\tilde{\alpha}\|$. That is: the maximum bias is the covariance of the approximation errors of the short representations, which is upper bounded by the product of the $L^{2}$ errors. The implication of this result is that there is a sharp bias bound that doesn’t depend on the full complexity of the (“long”) data-generating distribution, only on how well the short regressions approximate the long regressions in mean-squared error.
We are able to extend these findings to the hybrid control trial scenario in which we use the trial-specific average treatment effect, which is also a partially linear functional of the conditional expectation function of the outcome. The following fits the trial-specific ATE into the methodology proposed by Chernozhukov et al. [14].
@@ 3.2.2 Deriving the Riesz Representer
$\alpha(A,X,D)=\frac{1}{q}\left(\frac{DA}{p(X)}-\frac{D(1-A)+(1-D){r}(X)}{{\pi}(X)(1-{p}(X))+(1-{\pi}(X)){r}(X)}\right)\pi(X).$
@@ 3.2.3 Measuring Unobserved Confounding
The methodology proposed by Chernozhukov et al. [14] relies on the outcome regression and the Riesz representer, which have been identified above, and gains in explanatory power from latent, or unobserved, confounders. We can formally define $C^{2}_{Y}$ and $C^{2}_{D}$ as the gain in information from including unobserved confounders $Z$ in the outcome regression and Riesz representer in terms of the non-parametric $R^{2}$, letting the tilde denote the long regression in which the unobserved confounders are included in estimation of the functional components.
$C^{2}_{D}$ $=\frac{E[\tilde{\alpha}(A,X,D)^{2}]-E[\alpha(A,X,D)^{2}]}{E[\alpha(A,X,D)^{2}]}$
$C^{2}_{Y}$ $=\frac{E[(\tilde{\mu}(A,X)-\mu(A,X))^{2}]}{E[(Y-\mu(A,X)^{2}]}$
@@ 3.2.4 Calculating the Bias Bound $B$
As in Chernozhukov et al. [14], we reformulate the squared-bias as
@@
$\rho$ $=Cor(\mu-\tilde{\mu},\alpha-\tilde{\alpha})$
where $\rho$ measures the degree to which confounding in $m$ is related to confounding in $\alpha$.
The quantity $S$ depends only the observable (“short”) distribution, whereas the other parameters depend on the long distribution. Therefore only $S$ can be estimated from the data. $\rho$, $C_{D}$, and $C_{Y}$ must be assumed or bounded on the basis of domain knowledge. There are, however, sensible ways of setting these parameters. It is possible to pick a good value for $\rho$ — as well as $C^{2}_{D}$ and $C^{2}_{Y}$ — empirically through benchmarking methods using the observed covariates. For example, given a dataset with multiple observed covariates, one can select an observed covariate to be treated as unobserved and calculate estimates for $m$, $\tilde{m}$, $\alpha$, and $\tilde{\alpha}$ to then use to estimate $\rho$, $C^{2}_{D}$, and $C^{2}_{Y}$ through a “leave-one-out” approach [32]. The decision of which variable to treat as unobserved depends on multiple factors such as the nature of confounding we might expect to be present as well as the potential magnitude of confounding. However, there is no free lunch: any method for choosing these parameters incurs some risk and therefore requires a good understanding of the problem domain. In practice these parameters should also be varied over plausible ranges — see Section 3.4 on critical causal gaps below.
@@
Under adversarial confounding when $\rho^{2}=1$, the bias bound is sharp with $B=C_{Y}C_{D}S$. In our simulations, we proceed with the assumption that $C_{D}$ and $C_{Y}$ have been well-chosen a priori (e.g. using one of the methods described above) since we want to separately evaluate the performance of our method from a user’s knowledge or assumptions. Certainly, our method will appear to fail if these parameters are poorly chosen.
3.2.5 EEE Estimator of $B$
In order to estimate $B$, we must first estimate $S$, which depends on our Riesz representer and outcome regressions. For the outcome regression $\mu(A,X)$, we estimate separate regressions for each treatment. To estimate $\mu(1,X)$, we use only the data corresponding to those who were part of the original RCT, regressing $Y$ onto $A$ and $X$ for only $D=1$. However, due to the mean exchangeability assumption we can estimate $\mu(0,X)$ with both the RCT and external data arms (thus where we gain efficiency in the estimation of the trial-specific ATE), regressing $Y$ onto $A$ and $X$ regardless of $D$. Although pooling the data to estimate $\mu(0,X)$ when mean exchangeability does not hold introduces bias, this is precisely the bias that we are measuring in this paper. If we can discern the magnitude of the bias, we can evaluate whether the risk of bias from pooling outweighs the benefit of increased efficiency from pooling.
@@ 3.2.5 EEE Estimator of $B$
3.2.6 Inference on $B$
Since the bias bound requires estimation of the term $S$, it is itself a random quantity with some uncertainty that should be quantified. Inference on $B$ (via $S$) can be conducted using the delta method to combine estimates of the standard errors of $\nu^{2}=E[\alpha(X,A,D)]^{2}$ and $\sigma^{2}=E[Y-\mu(A,X)]^{2}$. The influence functions for $\nu^{2}$ and $\sigma^{2}$ are $\phi_{\sigma^{2}}=(Y-\mu(A,X))^{2}-\sigma^{2}$ and $\phi_{\nu^{2}}=\alpha(X,A,D)^{2}-\nu^{2}$. See Chernozhukov et al. [14] for additional details. The influence function of $B$ is then:
@@ 3.3 Using the Bias Bound
Presuming that we had a true upper bound $B\geq b$ and the true value of the observable trial-specific ATE $\Psi$, we could define a range $[\Psi_{-},\Psi_{+}]$ (calculated using $\Psi_{\pm}=\Psi\pm B$) which would be certain to include the true, causal, trial-specific ATE. Our task is therefore to produce estimates and conduct inference for the estimands $\Psi_{\pm}$ using our efficient estimates of $\Psi$ and $B$ which are described above.
3.3.1 Inference on $[\Psi_{-},\Psi_{+}]$
@@
Instead of constructing a two-sided confidence interval around the estimates of the bounds (such as one would do for inference on $\hat{\Psi}$), we follow Imbens and Manski [22] to construct one-sided confidence bounds on $\Psi_{+}$ and $\Psi_{-}$ that maintain uniform asymptotic coverage across $B$.
@@
[…]
Given a critical value of $\eta$, this confidence region that uses both confidence bounds guarantees coverage of $\Psi$ to asymptotically approach $1-\eta$ uniformly in $B$. This is important when there is very little unmeasured confounding. Another benefit of this procedure is that whenever $B\approx 0$, the confidence region becomes equivalent to that of a naive HCT analysis with a standard two-sided pointwise confidence interval. More details and justification can be found in Imbens and Manski [22].

### v2

@@ 2.2.1 Efficient Estimation
We rely on an efficient estimator $\hat{\Psi}$ for the trial-specific average treatment effect proposed by li_improving_2023. Although others exist, we focus on this estimator because it is asymptotically efficient relative to regular and asymptotically linear alternatives. This property aligns with our goal to alleviate trial constraints since efficient estimators permit smaller sample sizes while preserving standard methods of inference and robust behavior across data-generating processes [laan_targeted_2011]. This estimator is also doubly robust, making it less prone to model misspecification bias.
The efficient influence function for this estimand (as derived by li_improving_2023) is:
@@
In our setting, this is the equivalent of the popular augmented inverse propensity weighted (AIPW) estimator that is used in observational studies. It is also (when calculated with sample-splitting) the double machine learning (DML) or “efficient estimating equations” estimator for our setting [chernozhukov2024doubledebiasedmachinelearningtreatment]. Although we use DML here for accessibility and consistency with chernozhukov_long_2024, the work most closely related to ours, other (asymptotically equivalent) approaches for efficient estimation such as targeted maximum likelihood estimation (TMLE) can also be used [laan_targeted_2011].
@@
When one assumes mean exchangeability and has a binarized outcome $Y$, $r(X)$ reduces to 1. In this paper, we will restrict the outcome to be binary so that $r(X)$ does not require estimation and the results can be presented more economically. However, the results should still extend well to other cases (e.g. continuous or even censored outcomes). For example, the value of $r(X)$ only affects efficiency and not unbiased estimation, and $r(X)$ can be estimated well in a scenario where a continuous outcome is of interest as shown in li_improving_2023.
@@ 3 Methods
The issue at hand is that this estimator, like others, can only target the observable estimand which relies on the strict mean exchangeability assumption in order to get any benefit from the external control subjects. Although there are approaches such as the test-then-pool method to detect violations of this assumption, they cannot recognize every violation [viele_use_2014]. Further, while preliminary tests may identify a violation of the mean exchangeability assumption, it’s possible that the violations only lead to small amounts of bias. In these cases, the external data may have been useful, providing greater benefits than costs, but not used due to the black-and-white nature of significance tests. More complicated methods of dynamic borrowing include Bayesian methods, but simulation studies have shown that power gains are limited when strict type-one error rates are enforced [koppschneider_power_2020]. Due to the inevitability and uncertain quantification of this bias, hybrid control trials are typically not accepted by major regulatory bodies [noauthor_international_2000].
@@ 3.1 Omitted Variable Bias
[…]
In what follows, our goal will be to find an upper-bound $|\Psi-\Psi^{*}|\leq B$, quantifying the maximum bias that could be introduced if the mean exchangeability assumption were violated. We will then derive an estimator $\hat{B}$ for $B$, the bias bound.
@@ 3.2 The Bias Bound $B$
We leverage prior omitted variable bias methodologies for developing sensitivity analyses based on the Riesz representation form of the causal estimand [chernozhukov_long_2024]. Prior sensitivity analyses have been developed for a variety of estimands in observational studies such as the average treatment effect and weighted average incremental effects. In particular, this method works for a class estimands that can be written as a linear functional of the conditional expectation function of the outcome, also known as the outcome regression. These estimands can be written in the form: $\Psi(P_{0})=E[m(\mu)(O)]$ where $\mu$ is the outcome regression based on $P_{0}$ and $m$ is some linear functional.
@@ 3.2.1 Riesz Representation
The Riesz representation theorem states that within a Hilbert space any continuous linear function that maps to a real number can be written as the inner product of the argument to the function and some other element in the Hilbert space [intro, lee2025rieszboostgradientboostingriesz]. Using this theorem, we can then rewrite estimands of the form given above as $E[m(\mu)(O)]=E[\mu(O)\alpha(O)]$ where $\alpha$ is defined as the Riesz representer.
As an example, we can consider the well-known estimand for the average treatment effect of a binary treatment: $\Psi(P_{0})=E[E[Y|A=1,X]-E[Y|A=0,X]]$. The ATE can be defined as a linear functional of the outcome regression with $m(\mu)(O)=\mu(1,X)-\mu(0,X)$. Standard conditioning arguments show that the Riesz representer for the ATE is $\frac{A}{P(A=1|X)}-\frac{1-A}{P(A=0|X)}$, the usual inverse propensity weights. In this way, the Riesz representation motivates the use of such weighted estimators [lee2025rieszboostgradientboostingriesz].
chernozhukov_long_2024 uses this representation to define a bias bound $B$ for any Riesz-representable estimand, dependent on the residuals between the long and short formulations of the outcome regression and Riesz representer. Since the short regressions are projections of the long regressions onto a smaller space of functions depending only on $X$, these short regressions are orthogonal to the additional regression components living outside of this smaller subspace. Relying on the orthogonality of $\mu$ and $\alpha$ on these residuals, it is shown that the bias (in terms of the difference of the long and short estimands) is bounded: $\Psi-\tilde{\Psi}=E[(\mu(O)-\tilde{\mu}(O))(\alpha(O)-\tilde{\alpha}(O))]\leq\|\mu-\tilde{\mu}\|\|\alpha-\tilde{\alpha}\|$. That is: the maximum bias is the covariance of the approximation errors of the short representations, which is upper bounded by the product of the $L^{2}$ errors. The implication of this result is that there is a sharp bias bound that doesn’t depend on the full complexity of the (“long”) data-generating distribution, only on how well the short regressions approximate the long regressions in mean-squared error. Another nice implication of this result is that we can concern ourselves only with confounders that affect both the outcome regression $\mu$ and the Riesz Representer $\alpha$, given the product term in the expectation. If a confounder only affects one of the terms, say $\mu$, the bias will be equal to zero since $\alpha(O)-\tilde{\alpha}(O)$ will also be equal to zero. This implication will later be further discussed in the context of HCTs.
We are able to extend these findings to the hybrid control trial scenario in which we use the trial-specific average treatment effect, which is a partially linear functional of the conditional expectation function of the outcome. The following fits the trial-specific ATE into the methodology proposed by chernozhukov_long_2024.
@@ 3.2.2 Deriving the Riesz Representer
$\alpha(A,X,D)=\frac{1}{q}\left(\frac{DA}{p(X)}-\frac{D(1-A)+(1-D){r}(X)}{{\pi}(X)(1-{p}(X))+(1-{\pi}(X)){r}(X)}\pi(X)\right).$
@@ 3.2.3 Measuring Unobserved Confounding
The methodology proposed by chernozhukov_long_2024 relies on the outcome regression and the Riesz representer, which have been identified above, and gains in explanatory power from latent, or unobserved, confounders. We can formally define $C^{2}_{Y}$ and $C^{2}_{D}$ as the gain in information from including unobserved confounders $Z$ in the outcome regression and Riesz representer in terms of the non-parametric $R^{2}$. We let the tilde denote the long regression in which $\tilde{X}$, the set of observed covariates and unobserved confounders, is used in the estimation of the functional components. The functional components of the Riesz Representer in this case are $r(X)$, $p(X)$, and $\pi(X)$. Since $r(X)$ reduces to 1 and $p(X)$ is roughly set to some value in an RCT, the primary gains in explanatory power in the Riesz Representer likely come from estimation of $\pi(X)$. We can thus think of the gains in explanatory power of the Riesz Representer primarily focusing on the gains in explanatory power a confounder may have on the trial selection propensity $\pi(X)$ – note that this is not fully the case but we believe that this may be helpful for understanding what $C^{2}_{D}$ measures. In other words, $C^{2}_{D}$ roughly measures how well the observed covariates capture the differences between the external and RCT populations.
$C^{2}_{D}$ $=\frac{E[\tilde{\alpha}(A,\tilde{X},D)^{2}]-E[\alpha(A,X,D)^{2}]}{E[\alpha(A,X,D)^{2}]}$
$C^{2}_{Y}$ $=\frac{E[(\tilde{\mu}(A,\tilde{X})-\mu(A,X))^{2}]}{E[(Y-\mu(A,X))^{2}]}$
Note that the non-parametric partial $R^{2}$ is not bounded similarly to a typical measure of $R^{2}$. The gains in explanatory power are relative to what is already explained by the set of observed covariates $X$; if the gains from including confounders exceed what is already explained, the numerator of the equations for $C^{2}_{Y}$ and $C^{2}_{D}$ will be larger than the denominators. Therefore, it is possible for these values to exceed 1 unlike a typical measure of $R^{2}$. Further, since the measure is relative to what is already captured by the observed covariates $X$, it is possible that some confounder $Z$ has a strong relationship with the outcome or Riesz Representer but a small non-parametric $R^{2}$ if a covariate exists in $X$ that already captures some mechanism of $Z$ (ex. a confounder such as diet may have a strong relationship with some outcome but part of the relationship may be captured by an observed covariate such as BMI or cholesterol levels).
@@ 3.2.4 Calculating the Bias Bound $B$
As in chernozhukov_long_2024, we reformulate the squared-bias as
@@
[…]
where $\rho$ measures the degree to which confounding in $\mu$ is related to confounding in $\alpha$.
The quantity $S$ depends only the observable (“short”) distribution, whereas the other parameters depend on the long distribution. Therefore only $S$ can be estimated from the data. $\rho$, $C^{2}_{Y}$, and $C^{2}_{D}$ must be assumed or bounded on the basis of domain knowledge, thus we will refer to these as the sensitivity parameters. There are, however, sensible ways of setting these parameters. It is possible to pick a good value for $\rho$ — as well as $C^{2}_{D}$ and $C^{2}_{Y}$ — empirically through benchmarking methods using the observed covariates. For example, given a dataset with multiple observed covariates, one can select an observed covariate to be treated as unobserved and calculate estimates for $m$, $\tilde{m}$, $\alpha$, and $\tilde{\alpha}$ to then use to estimate $\rho$, $C^{2}_{D}$, and $C^{2}_{Y}$ through a “leave-one-out” approach [mcclean_calibrated_2024]. The decision of which variable to treat as unobserved depends on multiple factors such as the nature of confounding we might expect to be present as well as the potential magnitude of confounding. However, there is no free lunch: any method for choosing these parameters incurs some risk and therefore requires a good understanding of the problem domain. In practice these parameters should also be varied over plausible ranges — see Section 3.4 on critical causal gaps below.
Finally, also note that the bias bound relies on products of the sensitivity parameters. Therefore, if any of these are 0, the resulting bias from confounding is also 0. Similarly, if any one of the parameters is large but the others are ‘small’, the resulting bias most likely will also be relatively small. In the context of HCTs, this translates to confounding must jointly affect (measured by $\rho$) both the outcome regression and Riesz Representer. Therefore, confounders must affect both trial participation and the outcome. Practically speaking, if external data is chosen well, we can somewhat control the magnitude of $C^{2}_{D}$ and $B$. In theory, even if there exists a confounder which greatly influences the outcome, if the confounder does not explain differences in the trial and external control populations, bias will not be incurred. Similarly, if there exists a confounder which greatly influences the Riesz Representer/trial participation but not the outcome, bias will not be incurred – though this is less controllable through selection of external controls, but may be useful in the discussion of possible confounders.
@@
Under adversarial confounding when $\rho^{2}=1$, the bias bound is sharp with $B=C_{Y}C_{D}S$. In our simulations, we proceed with the assumption that $C_{D}$ and $C_{Y}$ have been well-chosen a-priori (e.g. using one of the methods described above) since we want to separately evaluate the performance of our method from a user’s knowledge or assumptions. Certainly, our method will appear to fail if these parameters are poorly chosen.
3.2.5 Estimator of the bias bound
In order to estimate $B$, we must first estimate $S$, which depends on our Riesz representer and outcome regressions. For the outcome regression $\mu(A,X)$, we estimate separate regressions for each treatment. To estimate $\mu(1,X)$, we use only the data corresponding to those who received treatment, regressing $Y$ onto $X$ for only $A=1$ (note this implied $D=1$). However, under the mean exchangeability assumption, we estimate $\mu(0,X)$ using both the RCT and external control arms (this is where we gain efficiency in the estimation of the trial-specific ATE). This corresponds to regressing $Y$ on $X$ for $A=0$, regardless of the value of $D$. Although such pooling introduces bias when mean exchangeability fails (since $\mu(0,X)$ is no longer identified by $E[Y|A=0,X]$), this is precisely the bias that we are measuring in this paper. If we can discern the magnitude of the bias, we can evaluate whether the potential bias from pooling outweighs the efficiency gains it provides.
@@ 3.2.5 Estimator of the bias bound
3.2.6 Inference on the bias bound
Since the bias bound requires estimation of the term $S$, it is itself a random quantity with some uncertainty that should be quantified. Inference on $B$ (via $S$) can be conducted using the delta method to combine estimates of the standard errors of $\nu^{2}=E[\alpha(X,A,D)]^{2}$ and $\sigma^{2}=E[Y-\mu(A,X)]^{2}$. The influence functions for $\nu^{2}$ and $\sigma^{2}$ are $\phi_{\sigma^{2}}=(Y-\mu(A,X))^{2}-\sigma^{2}$ and $\phi_{\nu^{2}}=\alpha(X,A,D)^{2}-\nu^{2}$. See chernozhukov_long_2024 for additional details. The influence function of $B$ is then:
@@ 3.3 Using the Bias Bound
Presuming that we had a true upper bound $B\geq b$ and the true value of the observable trial-specific ATE $\Psi$, we could define a range $\Psi_{-},\Psi_{+}]$ (calculated using $\Psi_{\pm}=\Psi\pm B$) which would be certain to include the true, causal, trial-specific ATE. Our task is therefore to produce estimates and conduct inference for the estimands $\Psi_{\pm}$ using our efficient estimates of $\Psi$ and $B$ which are described above.
3.3.1 Inference on $\Psi_{-},\Psi_{+}]$
@@
Instead of constructing a two-sided confidence interval around the estimates of the bounds (such as one would do for inference on $\hat{\Psi}$), we follow imbens to construct one-sided confidence bounds on $\Psi_{+}$ and $\Psi_{-}$ that maintain uniform asymptotic coverage across $B$.
@@
where $\Phi(c)=\int_{-\infty}^{c}\frac{1}{2\pi}e^{\frac{-x^{2}}{2}}dx$.
Given a critical value of $\eta$, this confidence region that uses both confidence bounds guarantees coverage of $\Psi$ to asymptotically approach $1-\eta$ uniformly in $B$. This is important when there is very little unmeasured confounding. Another benefit of this procedure is that whenever $B\approx 0$, the confidence region becomes equivalent to that of a naive HCT analysis with a standard two-sided pointwise confidence interval. More details and justification can be found in imbens.